# HyperFusion: A Hypernetwork Approach to Multimodal Integration of Tabular and Medical Imaging Data for Predictive Modeling

### v1

@@ 3.1.1 Embedding the tabular data
For our tabular embedding network, we opted for an MLP model given that tabular data lacks spatial context. This choice aligns with its favorable performance and suitability for small datasets, as noted in (Borisov et al., 2022). However, the hypernetwork configuration remains versatile, allowing the integration of any desired embedding architecture.
The proposed framework is end-to-end and the embedding parameters are learned concurrently with the entire network, as illustrated in Figures 1 and 2B. Moreover, the continuity of the embedding space, where similar samples are mapped closer together, enables interpolation. This capability allows the model to generate meaningful representations for previously unseen samples, enhancing its ability for generalization
@@ 3.1.2 Weights initialization
A network’s convergence is contingent upon the initial distribution of its parameters. Typically, neural network layer parameters are initialized with consideration to their fan-in or fan-out, a practice that stabilizes their convergence, as outlined in (He et al., 2015a; Glorot and Bengio, 2010). In our context, parameter initialization assumes a crucial role, particularly concerning $\theta_{\mathcal{H}}$ generated by the hypernetwork (its output) for the primary network. These parameters are computed as part of the hypernetwork’s feedforward process and cannot be directly initialized to follow a specific distribution. Furthermore, these parameters vary with each different input sample $T$.
To address this challenge, we employed the technique detailed in (Chang et al., 2019), which conducted variance analysis on the hypernetwork’s output, $\theta_{\mathcal{H}}$. The method aims to initialize the internal hypernetwork parameters ($\phi$) in a way that enables the feedforward process of the hypernetwork to generate external primary network parameters ($\theta_{h_{i}}$) fitting a specific distribution. More precisely, the distribution of these parameters should ensure that the output variance in their designated layers within the primary network closely aligns with the input variance. This initialization contributes to the convergence of the proposed hypernetwork framework.
@@ 3.2 Loss functions
As depicted in Figure 1 the loss backpropagates throughout the entire network compound affecting both the internal primary network’s parameters and the hypernetwork. Given the flexibility of our framework, the loss function can be adapted to a wide range of applications. It consists of two components the task-specific loss, $\mathcal{L}_{\mbox{\footnotesize{task}}}$, and a weight decay regularization term $\mathcal{L}_{\mbox{\footnotesize regularization}}$ as follows:
@@
where, $y$ are the true labels/values. The regularization is applied to $\phi$ and $\theta_{\mathcal{P}}$ to reduce overfitting. Recall that the external primary network parameters produced by the hypernetwork, i.e., $\theta_{\mathcal{H}}$ do not go through the gradient descent process and therefore are not regularized.
@@
For classification tasks with $C$ classes, aiming to account for potential class imbalance, we use the Weighted Cross-Entropy (WCE) loss. The weight of each class $c$ is denoted as $w^{c}$, and is inversely proportional to the frequency of the class in the training and validation sets. The WCE loss is formulated as follows:
@@ 3.3 Missing values
It is not uncommon to have tabular data with missing values. Handling missing values is a crucial aspect of tabular data preprocessing, as it can significantly impact the quality of modeling and analysis. There exist different strategies to address this challenge. For example, the multi-modal generative approach proposed in (Dolci et al., 2022) which compensates for missing modalities via a three-module framework. In the first module, each modality is processed independently, the second one imputes the unavailable data using pretrained generators and the last module fuses all features prior to the prediction.
In our study, we impute missing values using an iterative approach. The tabular data is presented as a matrix where each column represents an attribute and each row is a training set sample. In each step, one attribute (column) is selected as the output $z$, and the remaining columns are considered as inputs $X$. A regressor is trained on the non missing values of $(z,X)$, and then employed to predict the missing values of $z$. This iterative process repeats for each of the columns. As in (Wolf et al., 2022) we add a indicator, such as a NaN flag to indicate imputed values.
@@ 3.4 Ensemble learning
In multiclass classification tasks involving $C$ classes, each model $\mathcal{F}_{m}$ produces a probability distribution vector over the predicted classes, represented as $\mathbf{p}_{m}=(p^{c=1},\ldots,p^{c=C})_{m}$. To obtain the final ensemble prediction, $\mathcal{E}_{classification}(T,I)$, these models are combined using the weighted average of their probability distributions, as follows:
@@ 3.5.1 Conditioned brain age prediction using hypernetworks
The proposed hypernetwork framework for conditioned brain age prediction is illustrated in Figure 3. The primary network takes 3D brain MRI scans (T1w) of healthy subjects as input, while the tabular data fed into the hypernetwork consists of binary values indicating the subjects’ sex.
The hypernetwork compound is trained to address a regression problem using the loss function defined in Equations 1 and 2. The primary network architecture (Figure 3B-C) is built upon the VGG backbone (Simonyan and Zisserman, 2014), as suggested in (Levakov et al., 2020) for brain age prediction.
The hypernetwork itself consists of a simple linear layer that embeds the incoming binary tabular data through a 2-to-1 layer configuration, as depicted in Figure 3D. The embedding step is crucial, especially when using the orthogonal one-hot representation of the sex attribute (01 or 10). Without embedding, the hypernetwork would generate a distinct set of weights for each sex attribute value. We found that in this case, the most effective integration of the hypernetwork occurred within the linear layers at the final processing stage.
Figure 3: The proposed HyperFusion architecture for the conditioned brain age prediction task. A: The inputs include the subjects’ sex (encoded as a 2D one-hot vector) representing a single attribute tabular data and the corresponding 3D brain MRIs. B: The primary network backbone is a variant of the VGG architecture(Simonyan and Zisserman, 2014), where the parameters of its final four linear layers (framed in red) are external and generated by the hypernetwork. The network output the subjects’ brain age ($\in\mathbb{R}$), and the loss function is a weighted sum of the regression loss (MSE) and weight decay regularization, as defined in Equations 1-2. C: A closer look of a convolutional block, comprising two convolutional layers with ReLU activation, batch normalization and downsampling through max-pooling. D: The hypernetwork comprises four sub-networks ($h_{1},h_{2},h_{3},h_{4}$), each corresponds to one of the linear layer in the primary network. Each $h_{k}$ is composed of a 2-to-1 MLP for the embedding of the one-hot input vectors.
@@ 3.5.2 AD classification using hypernetworks
Figure 4 visually presents the hypernetwork framework for multi-class classification of subjects into CN, MCI, and AD groups. The input to the hypernetwork (Figure  4A) is the tabular data which comprise nine clinical and demographic attributes as described in Section 4.2.3. Similar to the brain age prediction’s application, the primary network is fed with 3D brain MRI scans (Figure  4B). Yet, for AD classification, the images were cropped to include only the hippocampus and its surrounding tissues as suggested in (Wen et al., 2020). The cropping significantly reduced the input size leading eventually to a reduction in the number of the required network parameters and to better classification results. We note that the hippocampus is the significant component of the limbic lobe being a crucial region for learning and memory. Hippocampal atrophy is a known bio-marker for AD (Rao et al., 2022; Salta et al., 2023). The hypernetwork, illustrated in Figure 4C is composed of a non-linear, single-layer MLP used for embedding, with Parametric ReLU (P-ReLU) (He et al., 2015b). The MLP generates weights and biases for the primary network. The primary network predictions are probability distributions ($P_{CN},P_{MCI},P_{AD}$) produced by the softmax layer. The loss is a weighted sum of the multi-class classification loss (WCE) and a weight decay regularization term, as detailed in Equations 1,3. The backbone of the primary network, depicted in Figure 4B is based on the pre-activation ResNet blocks followed by two linear layers. The last ResNet block (framed in red), called the HyperRes block, gets a subset of its parameters from the hypernetwork. A closer view of a pre-activation ResNet blocks is presented in Figure 4D. The architecture (based on (He et al., 2016)) is the same for all blocks, yet the illustration specifically refers to the HyperRes block which gets the parameters to one of its convolutional layer (framed in red) from the hypernetwork.
Figure 4: The proposed HyperFusion architecture for the AD classification task. A: The input to the network compound include clinical and demographic tabular attributes ($d$ in total) of the subjects along with their brain MRIs. B: The primary network’s is composed of pre-activation ResNet blocks followed by two linear layers. The last ResNet block (HyperRes block, framed in red) gets a subset of its parameters from the hypernetwork. The primary network predictions are probability distributions ($P_{CN},P_{MCI},P_{AD}$) produced by the softmax layer. The loss is a weighted sum of the classification loss (weighted CE) and weight decay regularization, as detailed in Equations 1,3. C: The hypernetwork is composed of an MLP for embedding, with Parametric ReLU (P-ReLU) (He et al., 2015b). The MLP generates weights and biases for the primary network. D: A closer look at the pre-activation Res Block. The architecture (based on (He et al., 2016)) is the same for all blocks, yet the illustration specifically refers to the HyperRes block which gets the parameters to its convolutional layer (framed in red) from the hypernetwork.
@@ 4 Experiments
In this section, we elucidate the experimental configurations, ablation studies, prediction outcomes, and comparisons for two diverse brain imaging analysis applications, aiming to showcase the robustness and adaptability of the proposed multi-modal fusion hypernetwork framework. In particular, Section 4.1 delves into the experiments and findings related to brain age estimation conditioned by the subject’s sex, while Section 4.2 addresses multi-class classification of subjects into CN, MCI, or AD groups.
@@ 4.1.1 The data
The data for this study include 26,691 brain MRI scans of different healthy human subjects from 19 sources. The table in Appendix A presents a comprehensive information on each dataset including the number of subjects as well as their age and sex distributions.
@@ 4.1.2 Preprocessing and partitioning
We applied a preprocessing pipeline to the T1-weighted MRI scans as proposed in (Levakov et al., 2020) - see Appendix A for details. Before being utilized as model inputs, the images were standardized to attain a mean of zero and a standard deviation of one across all non-zero voxels, ensuring data uniformity for an effective training. We note that subjects’ sex and age are routinely recorded during the scanning visits. Therefore we can utilize almost all of the available scans without having to handle issues related to missing values. The tabular data, in this case, consists of a single binary value which is the subject’s sex - male or female. It is represented in out study by a 2D one-hot vector. We partitioned the dataset into training (80%), validation (10%), and test (10%) subsets, preserving consistent age and sex distributions across all the subsets.
@@ 4.1.3 Differences in brain aging between males and females
The results of this experiment assessed our hypothesis (Figure 5A): Models that were trained on single sex data achieved the highest performance on test data acquired from the people of the corresponding sex (among all three models) Confirming the validity of our hypothesis enables us to proceed and incorporate the hypernetwork.
Figure 5: Conditioned brain age prediction results. A) An imaging only experiment. A presentation of the MAE for baseline networks trained on either male only (blue), female only (pink) and mixed (green) equally sized training sets, where the test sets are composed of male only (left), female only (middle) or mixed (right) subsets. The plot demonstrate that when a network is trained and tested on a same-sex population the MAE results are enhanced, assessing the hypothesis that aging processes are influenced by the subjects’ sex. B) HyperFusion: Images and Sex. A comparison between the baseline model (orange) using imaging data alone and the proposed HyperFusion model (purple) exploiting both imaging data and sex information in terms of MAE. The test was performed using either males subset (left), females subset (middle) or the full dataset (right).
4.1.4 Training and evaluation
We used the Mean Average Error (MAE) between the chronological and the predicted age for evaluating our framework. To ensure robustness, we trained multiple models with different random weight initialization mitigating the potential impact of incidental results. The test results were calculated based on an ensemble of the five highest-performing models identified during validation. The ensemble’s output was derived through an average of the models predictions, as detailed in Section 3.4, further enhancing the reliability of our findings.

### v2

@@ 3.1.1 Embedding the tabular data
For our tabular embedding network, we opted for an MLP model given that tabular data lacks spatial context. This choice aligns with its favorable performance and suitability for small datasets, as noted in [5]. However, the hypernetwork configuration remains versatile, allowing the integration of any desired embedding architecture.
The proposed framework is end-to-end and the embedding parameters are learned concurrently with the entire network, as illustrated in Figure 1. Moreover, the continuity of the embedding space, where similar samples are mapped closer together, enables interpolation. This capability allows the model to generate meaningful representations for previously unseen samples, enhancing its ability for generalization
3.1.2 Hyperlayer’s position selection
While the proposed hypernetwork concept is general, its adaptation to different applications requires the accommodation to specific architectures of the primary network. Inspired by the layer selection method proposed in  [39], we identify ‘good candidates’ for hyperlayers by evaluating the impact of each primary network layer given the untrained, backbone network. The evaluation is carried out by randomly initializing the parameters of the selected layer multiple times ($N=1000$) while keeping the random parameters of the other network layers fixed. The entropy of the layer-based loss is calculated based on the normalized histogram of the loss values obtained for each random initialization of the layer of interest. Final layer selection from the pool of candidates is conducted empirically as demonstrated in the ablation study in D.
@@ 3.1.3 Weights initialization
A network’s convergence is contingent upon the initial distribution of its parameters. Typically, neural network layer parameters are initialized taking into account their fan-in or fan-out, a practice that stabilizes their convergence, as outlined in [25, 21]. In our context, parameter initialization assumes a crucial role, particularly concerning $\theta_{\mathcal{H}}$ generated by the hypernetwork (its output) for the primary network. These parameters are computed as part of the hypernetwork’s feedforward process and cannot be directly initialized to follow a specific distribution. Furthermore, these parameters vary with each different input sample $T$.
To address this challenge, we employ the technique detailed in [8], which performed variance analysis on the hypernetwork output, $\theta_{\mathcal{H}}$. The method aims to initialize the parameters $\phi_{k}$ of each hypernetwork $k$ in a way that enables the hypernetwork feedforward process to generate external primary network parameters ($\theta_{h_{k}}$) fitting a specific distribution. More precisely, the distribution of these parameters should ensure that the output variance in their designated layers within the primary network is closely aligned with the input variance. This initialization contributes to the convergence of the proposed hypernetwork framework.
Formally, let $\phi_{k}=\{H^{W}_{k},H^{B}_{k}\}$ denote the parameters of the fully connected layers of $h_{k}$ that respectively generate the weights $W^{j}_{k}\in\theta_{h_{k}}$ and the biases $B^{j}_{k}\in\theta_{h_{k}}$ to a hyperlayer $j$ in the primary network. Let $d_{k}$ and $d_{j}$ denote the fan-in size of the hypernetwork and the primary network layers, respectively. We also denote by $e(T)$ the embedding (the output of the MLP) of the pre-processed tabular data and by $\mbox{Var}(\cdot)$ the variance. To achieve the desired variance, the parameters $H^{M}_{k}$ of each hypernetwork $k$ are initialized by sampling from a uniform distribution between $-\sqrt{3\mathcal{V}(H^{M}_{k})}$ and $\sqrt{3\mathcal{V}(H^{M}_{k})},$ where $M$ either stands for the weights $W$ or biases $B$ and $\mathcal{V}(H^{M}_{k})$ is calculated for each set of parameters, as in [8] as follows:
$\mathcal{V}(H^{W}_{k})=\frac{1}{d_{j}\cdot d_{k}\cdot\mbox{Var}(e(T))}$ (1)
$\mathcal{V}(H^{B}_{k})=\frac{1}{d_{k}\cdot\mbox{Var}(e(T))}$ (2)
@@ 3.2 Loss functions
As depicted by the yellow arrows in Figure 1 the loss backpropagates throughout the entire network compound affecting both the internal primary network’s parameters and the hypernetwork. Given the flexibility of our framework, the loss function can be adapted to a wide range of applications. It consists of two components the task-specific loss, $\mathcal{L}_{\mbox{\footnotesize{task}}}$, and a weight decay regularization term $\mathcal{L}_{\mbox{\footnotesize regularization}}$ as follows:
@@
where, $y$ are the true labels/values. Regularization is applied to $\phi$ and $\theta_{\mathcal{P}}$ to reduce overfitting. Recall that the external primary network parameters produced by the hypernetwork, i.e., $\theta_{\mathcal{H}}$ do not go through the gradient descent process and therefore are not regularized.
@@
For classification tasks with $C$ classes, in order to account for possible class imbalance, we use the Weighted Cross-Entropy (WCE) loss. The weight of each class $c$ is denoted as $w^{c}$, and is inversely proportional to the frequency of the class in the training and validation sets. The WCE loss is formulated as follows:
@@ 3.3 Missing values
It is not uncommon to have tabular data with missing values. Handling missing values is a crucial aspect of tabular data preprocessing, as it can significantly impact the quality of modeling and analysis. There exist different strategies to address this challenge. For example, the multi-modal generative approach proposed in [13] which compensates for missing modalities via a three-module framework. In the first module, each modality is processed independently, the second one imputes the unavailable data using pretrained generators and the last module fuses all features prior to the prediction.
In our study, we impute missing values using an iterative approach. The tabular data is presented as a matrix where each column represents an attribute and each row is a training set sample. In each step, one attribute (column) is selected as the output $z$, and the remaining columns are considered as inputs $X$. A regressor is trained on the non missing values of $(z,X)$, and then employed to predict the missing values of $z$. This iterative process repeats for each of the columns. As in [68] we add a indicator, such as a NaN flag to indicate imputed values.
@@ 3.4 Ensemble learning
In multi-class classification tasks involving $C$ classes, each model $\mathcal{F}_{m}$ produces a probability distribution vector over the predicted classes, represented as $\mathbf{p}_{m}=(p^{c=1},\ldots,p^{c=C})_{m}$. To obtain the final ensemble prediction, $\mathcal{E}_{classification}(T,I)$, these models are combined using the weighted average of their probability distributions, as follows:
@@ 3.5.1 Conditioned brain age prediction using hypernetworks
The proposed hypernetwork framework for conditioned brain age prediction is illustrated in Figure 2. The primary network takes 3D brain MRI scans (T1w) of healthy subjects as input, while the tabular data (single attribute) fed into the hypernetwork consists of 2D one-hot vectors representing the subjects’ sex. The hypernetwork compound is trained to address a regression problem using the loss function defined in Equations 3 and 4. The primary network architecture (Figure 2B-C) is a variant of the VGG backbone [60], as suggested in [35] for brain age prediction. The parameters of its final four linear layers (framed in red) are external and generated by the hypernetwork. The network output the subjects’ brain age ($\in\mathbb{R}$), and the loss function is a weighted sum of the regression loss (MSE) and weight decay regularization, as defined in Equations 3-4. Figure 2C provides a closer look of a convolutional block, comprising two convolutional layers with ReLU activation, batch normalization and downsampling through max-pooling. As illustrated in Figure 2D the hypernetwork comprises four sub-networks ($h_{1},h_{2},h_{3},h_{4}$), each corresponds to one of the linear layer in the primary network. Each $h_{k}$ is composed of a 2-to-1 MLP for the embedding of the one-hot input vectors. The embedding step is crucial, especially when using the orthogonal one-hot representation of the sex attribute (01 or 10). Without embedding, the hypernetwork would generate a distinct set of weights for each sex attribute value. We found that in this case, the most effective integration of the hypernetwork occurred within the linear layers at the final processing stage.
Figure 2: HyperFusion architecture for conditioned brain age prediction. A: The inputs include the subjects’ sex (encoded as a 2D one-hot vector) and the corresponding 3D brain MRIs. B: The primary network backbone is a variant of the VGG architecture [60], where the parameters of its final four linear layers (framed in red) are external and generated by the hypernetwork. C: A closer look of a convolutional block . D: The hypernetwork comprises four sub-networks ($h_{1},h_{2},h_{3},h_{4}$), each corresponds to one of the linear layer in the primary network.
@@ 3.5.2 AD classification using hypernetworks
Figure 3 visually presents the hypernetwork framework for multi-class classification of subjects into CN, MCI, and AD groups. The input to the hypernetwork (Figure  3A) is the tabular data which comprise nine clinical and demographic attributes as described in Section 4.3.3. Similar to the brain age prediction’s application, the primary network is fed with 3D brain MRI scans (Figure  3B). Yet, for AD classification, the images were cropped to include only the hippocampus and its surrounding tissues as suggested in [67]. The cropping significantly reduced the input size leading eventually to a reduction in the number of the required network parameters and to better classification results. We note that the hippocampus is the significant component of the limbic lobe being a crucial region for learning and memory. Hippocampal atrophy is a known bio-marker for AD [56, 58]. The hypernetwork, illustrated in Figure 3C is composed of a non-linear, single-layer MLP used for embedding, with Parametric ReLU (P-ReLU) [26]. The MLP generates weights and biases for the primary network. The primary network predictions are the probability distributions ($P_{CN},P_{MCI},P_{AD}$) produced by the softmax layer. The loss is a weighted sum of the multi-class classification loss (WCE) and a weight decay regularization term, as detailed in Equations 3,5. The backbone of the primary network, depicted in Figure 3B is based on the pre-activation ResNet blocks followed by two linear layers. The last ResNet block (framed in red), called the HyperRes block, gets a subset of its parameters from the hypernetwork. A closer view of a pre-activation ResNet blocks is presented in Figure 3D. The architecture (based on [27]) is the same for all blocks, yet the illustration specifically refers to the HyperRes block which gets the parameters to one of its convolutional layer (framed in red) from the hypernetwork.
Figure 3: HyperFusion architecture for the AD classification. A: The input consists of tabular attributes ($d$ in total) of the subjects along with their brain MRIs. B: The primary network’s is composed of pre-activation ResNet blocks followed by two linear layers. The last ResNet block (framed in red) gets a subset of its parameters from the hypernetwork. The primary network predictions are probability distributions ($P_{CN},P_{MCI},P_{AD}$) produced by the softmax layer. The loss is a weighted sum of the classification loss (weighted CE) and weight decay regularization, as detailed in Equations 3,5. C: The hypernetwork architecture D: A closer look at the pre-activation Res Block. See text for details.
@@ 4 Experiments
In this section, we detail the experimental configurations, ablation studies, prediction outcomes, and comparisons for two different brain imaging analysis applications, aiming to showcase the robustness and adaptability of the proposed multi-modal fusion hypernetwork framework. In particular, Section 4.2 delves into the experiments and findings related to brain age estimation conditioned by the subject’s sex, while Section 4.3 addresses multi-class classification of subjects into CN, MCI, or AD groups. Additional experiments for binary AD classification are presented in B.
4.1 Experimental details
All the experiments where conducted using a Tesla V100 32 GB GPU with PyTorch. Table 1 presents the experimental details of the AD classification and brain age prediction tasks. All architectural details of the networks are presented in Tables 56 in A.2 and C.2, respectively.
Table 1: Experimental details of the AD classification and brain age prediction tasks
Experiment epochs batch size optimizer weight decay factor learning rate
Brain age prediction 70 64 ADAM .05.05.05.05 $1.5\cdot 10^{-4}$
AD classification 250 32 ADAM $10^{-5}$ $10^{-4}$
@@ 4.2.1 The data
The data for this study include 26,691 brain MRI scans of different healthy human subjects from 19 sources. The table in A.1 presents a comprehensive information on each dataset including the number of subjects as well as their age and sex distributions.
4.2.2 Training and evaluation
We evaluate the hyperfusion and alternative brain-age prediction frameworks using the Mean Absolute Error (MAE) between the chronological and the predicted age. To ensure robustness, we trained multiple models with different random weight initialization to mitigate the potential impact of incidental results. The test results were calculated based on an ensemble of the five best-performing models identified during validation. The ensemble’s output was derived through an average of the model predictions, as detailed in Section 3.4, further enhancing the reliability of our findings.
@@ 4.2.3 Preprocessing and partitioning
We applied a preprocessing pipeline to the T1-weighted MRI scans as proposed in [35] - see  A.1 for details. Before being utilized as model inputs, the images were standardized to attain a mean of zero and a standard deviation of one across all non-zero voxels, ensuring data uniformity for effective training. We note that subjects’ sex and age are routinely recorded during the scanning visits. Therefore, we can utilize almost all of the available scans without having to handle issues related to missing values. The tabular data, in this case, consists of a single binary value which is the subject’s sex - male or female. It is represented in out study by a 2D one-hot vector. We partitioned the dataset into training (80%), validation (10%), and test (10%) subsets, preserving consistent age and sex distributions across all the subsets.
@@ 4.2.4 Differences in brain aging between males and females
The MAE scores of obtained by this set of experiments are presented in Figure 4A. The blue, pink and green bars refer to network training with either male only, female only or mixed data, respectively -maintaining equal-size training sets. The text sets are composed of male only (left), female only (middle) or mixed (right) subsets. The plot strengthens our hypothesis that sex information can facilitate brain age prediction - motivating the proposed hypernetwork framework.
Figure 4: Conditioned brain age prediction results. A) An imaging only experiment. The MAE scores obtained for baseline networks trained on either male only (blue), female only (pink) and mixed (green) equally sized training sets, where the test sets are composed of male only (left), female only (middle) or mixed (right) subsets. B) HyperFusion: A comparison between the baseline model using imaging data alone (orange), concatenation based fusion (light blue), Hyperfusion with different embedding - inputs 2d vector and outhputs 2d vector (olive green), and the proposed HyperFusion model (purple) using MAE metric. The inference was performed using either male test data (left), female test data (middle) or the entire (mixed) test dataset (right).
4.2.5 Comparisons and ablation study
Figure 4B presents comparisons of the proposed hyperfusion approach (purple bars) with both image-only (orange bars) and common tabular-imaging concatenation (light blue bars) methods. The comparisons were made for male-only (left bars), female-only (middle bars) and the entire (right bars) test data. For a fair comparison with our hyperfusion model, which provides parameters to all four linear layers in the primary network, the concatenation model combined the sex and image features in each linear layer of the baseline network. Additionally, we performed an ablation study on the embedding of the hypernetwork, specifically comparing our proposed one-hot vector into a single scalar mapping (purple bars) with a mapping into a 2D vector (olive-colored bars). An ablation study assessing the proposed external layer selection can be found in D.